MILLIMETER-WAVE SPECTROSCOPY OF THE HCN-H_{2} CLUSTER

\underline{M. ISHIGURO} and T. TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan; C. J. WHITHAM, K. HARADA and K. TANAKA, Institute for Molecular Science, Myodaiji, Okazaki 444-8585, Japan.

Millimeter-wave absorption spectroscopy combined with a pulsed-jet expansion technique was applied to the measurement of the rotational and rovibrational transitions of the HCN-H_{2} cluster in the frequency region of 75−150 GHz. So far, four rotational lines for the ground Σ_{0} state of the HCN-(o-H_{2}) cluster split into hyperfine structure due to the nitrogen nucleus were observed. Rotational constant B_{0} = 12899.718(20) MHz and centrifugal distortion constant D_{0} = 12.2470(16) MHz were derived together with its higher constants. The hyperfine constants determined eqQ = −2.830(33) MHz which is smaller than that of HCN molecule means a large amplitude motion of HCN of < θ > = 31.1 degree in the ground linear form. The bond length between HCN and H_{2} parts is derived to be 3.90 Å. Some lines belonging to the Σ_{1} − Σ_{0} van der Waals bending band of HCN-(o-H_{2}) were also observed. The Σ_{1} − Σ_{0} van der Waals bending frequency 136.831 GHz of HCN-(o-H_{2}) is larger than that of He-HCN 98.70 GHz, but much smaller than that of Ar-HCN 164.89 GHz. The hyperfine constant in the Σ_{1} state indicate the cluster has T-shape in the excited state of the vdW bending mode. A search of the rotational lines of the ground Σ_{0} state of HCN-(p-H2) and the Π_{0} state of HCN-(o-H_{2}) are now in progress as well as the vdW mode rovibrational lines for both HCN-(o-H_{2}) and -(p-H_{2}) clusters.